A sample of water has the following concentrations of ions (and pH = 7.0):

                        cations                                    mg/L      anions            mg/L

                        Ca+2      40.0                           HCO3-   110.0

                        Mg+2     10.0                           SO42-     67.2

                        Na+       ?                                 Cl-         11.0

                        K+         7.0

Assuming no other constituents are missing, use an anion-cation balance to estimate the concentration of Na+ (in mg/L as Na+)? Remember that the balance cannot be in mg/L.
Answer #1

(Ca2+) = (40 mg/L) x (meq/20 mg) x ( 50 mg CaCO3/meq) = 100 mg/L as CaCO3

(Mg2+) = (10 mg/L) x (meq/12.2 mg) x (50 mg CaCO3/meq) = 41 mg/L as CaCO3

(K+) = (7 mg/L) x (meq/39.1 mg) x (50 mg CaCO3/meq) = 9 mg/L as CaCO3

(Cl) = (11 mg/L) x (meq/35.5 mg) x (50 mg CaCO3 /meq) = 15.5 mg/L as CaCO3

(SO42-) = (67.2 mg/L) x (meq/48 mg) x (50 mg CaCO3 /meq) = 70 mg/L as CaCO3

(HCO3-) = (110 mg/L) x ( meq/61 mg) x (50 mg CaCO3 /meq) = 90.2 mg/L as CaCO3

(Na+) = \sum anions − \sum cations = 175.7- 150 = 25.7 mg/L as CaCO3

Therefore,

Na+ = (25.7 mg as CaCO3/L) x (23 mg/meq) x (meq/ 50 mg CaCO3) = 11.8 mg/L as CaCO3